# A comparative study of different TSO-DSO coordination in the reserve market

### v1

@@ Abstract
The increasing penetration of Distributed Energy Resources (DERs) in the distribution system has led to the emergence of a new market actor - the aggregator. The aggregator serves as a facilitator, enabling flexibility asset owners to get access to different markets. In which, EVs aggregators are gaining more attention due to their expanding use and potential to provide services in various types of markets, particularly in the reserve market. Currently, TSO indirectly utilizes these resources under the management of the distribution system operators (DSO), which can negatively impact the distribution grid. Conversely, adjustments from DSOs can impact service provision to TSO due to the shortage of TSO usage information. These factors highlight the importance of evaluating the service provision from aggregators under different TSO-DSO coordination schemes. This paper focuses on the provision of flexibility from electric vehicles (EVs) aggregators for balancing service in the TSO-DSO hybrid-managed and compares it with the DSO-managed coordination schemes. The behavior of aggregators reacting to price fluctuations and TSO requests under different coordination schemes and simulation scenarios is thoroughly evaluated. Additionally, their impact on the grid is analyzed through the DSO’s congestion management process and validated using data from a real part of the Dutch distribution network. Results find that the hybrid-managed coordination scheme gives more benefit to the aggregator than the DSO-managed scheme and the EVs aggregator will gain more profit in winter than summer due to more upward regulation service is needed.
@@
TSO-DSO coordination, DERs, aggregator, congestion, flexibility.
@@ I Introduction
Along with exploiting the potential of distributed energy resources, providing flexibility services with these sources attracts the attention of many organizations and research groups and has already been implemented in some European countries like France, Finland, Hungary, Estonia, Denmark, Belgium, Romania, etc [1]. The role of the aggregator - a market participant responsible for bundling flexibility from multiple small customers and producers into a portfolio and offering the combined capacity in other markets[2] - is becoming increasingly important. Many research studies focus on optimal biding strategies[3, 4, 5, 6, 7, 8], providing service to multiple markets (e.g., spot market, reserve,[3, 9, 10, 11, 12] congestion management services[7] or balancing portfolios[3]) and leveraging the potential of diverse appliances including electric vehicles (EVs), battery energy storage system (BESS), heat pumps, thermal-controlled load or their combination.
@@
Recently, EV aggregators are receiving increasing attention due to their expanding use, driven by environmental factors and government incentives. In 2023, the Netherlands was the sixth-largest EV market in Europe with more than 300 thousand battery electric vehicles registered and leads in EV infrastructure [13]. Research on the Dutch situation shows the potential of EV aggregators to provide service in the reserve market[10]. EV fleets also have the potential to provide flexibility to the internal balancing portfolio[3], and congestion management [7]. The above studies show that EV aggregators have great prospective to provide flexibility services and bring benefits not only for aggregators but also for prosumers and system operators. Therefore, the detailed EV aggregator is investigated in this research to consider its behaviors on different market price fluctuations.
@@
Moreover, the above research works focus on the aggregator business model when providing service to the transmission system operator (TSO) without a comprehensive analysis impact on the grid under the management of the distribution system operator (DSO). While an effective approach to leveraging DER flexibility is enhancing TSO-DSO coordination [14]. In recent years, TSO-DSO coordination has grown into a research trend of great interest to many researchers. Arthur et al [15] classify the existing coordination into three main schemes (1) TSO-managed; (2) DSO-managed; and (2) TSO-DSO Hybrid. The first one prioritizes the TSO in using flexibility, restricting the DSO from resources connected to their network. Conversely, the DSO-managed model assigns a greater role along with responsibility to DSO in optimizing the utility of flexibility. DSOs are prioritized to use the DER flexibility to manage congestion in their grids, then send the remainder to the central market. However, this scheme has the potential to create conflicts of benefit among DSOs. The Hybrid TSO-DSO managed model appears to be the most balanced option as TSO and DSO manage their own market and it allows TSO to use service with DSO’s consent. This coordination aligns well with the prevailing circumstances in many countries and offers distinct advantages over others by empowering DSOs role and fostering greater social welfare. Therefore, the DSO-TSO hybrid-managed model is one of the schemes selected for investigation in this research.
@@
Considering the provision of DERs flexibility from the aggregator within TSO-DSO coordination offers a broader perspective by highlighting the risk of violating grid constraints when procuring flexibility for the TSO. Therefore, this research constructs a comprehensive model for the EV aggregator to provide multiple services in different TSO-DSO coordination schemes. To the best of the authors’s knowledge, this topic has not been widely explored in the literature. This paper finds some similarities with the 3 papers [16, 17, 18] in using flexibility from the distribution grid via aggregator while considering the TSO-DSO coordination. However, unlike [16], this paper models the detail of EV aggregator, improves the DSO model with linearized ACOPF and compares the flexibility provision in 2 different coordination schemes. In [17], the authors show that 16% and nearly 52% of the operational cost is reduced for TSO and DSO, respectively. While aggregator lost about 54% profit in the DSO-managed model. They didn’t show how it would be different if TSO used flexibility from the aggregator in the balancing market. Compared with [18], this paper models the EV aggregator that considers individual EV to ensure EV energy within the limit range while providing vehicle to grid service and considering the TSO-DSO hybrid-managed model that advances over the DSO-managed coordination scheme.
@@
II TSO-DSO coordination with aggregator participation
Coordination between TSOs and DSOs can occur in various processes including pre-qualification, offering, and real-time activation of services. In the case of the Dutch balancing service, they coordinate pre-qualification of the equipment specifications before granting permission to connect to the grid. However, there is no interaction during the bidding and activation process, which takes place in real-time and near real-time. Balancing service is normally provided to TSO from the large generator and demand via a balancing service provider (BSP) [19]. A BSP offers a balancing service to TSO to balance out any unpredicted imbalances in the electricity grid. Two balancing services are used in the Netherlands, automatic frequency restoration reserve (aFRR) and manual frequency restoration reserve (mFRR). aFRR is frequently used to correct the imbalance within 15 minutes, while mFRR is used for larger and longer disturbances. Currently, BSP also takes the role of aggregator to provide these services by combining resources from scattered customers. In this study, coordination in the offering process for aFRR provision is proposed with two different schemes: the DSO-managed model and the TSO-DSO hybrid-managed model.
II-A TSO-DSO hybrid-managed
Figure 1: TSO-DSO hybrid-managed coordination scheme.
The first diagram is presented in Fig. 1, with 5 main processes. In which, the aggregator is responsible for aggregating all the resources from their prosumers, determining the optimal bidding strategy and sending it to the balancing responsible party (BRP) before the previous market gate closing date and determining the volume they will dedicate to service balancing. The TSO will make the economic dispatch offers and define the merit order list (MOL) (a list of offers sorted by price from low to high) for the next two imbalance settlement period (ISP), which is 30 minutes in the Netherlands. After receiving the MOL from the TSO, the DSO proceeds to validation, eliminate invalid bids/offers, and notify the TSO of the valid results. The processes are as follows:
  1. 1.
    Aggregator optimizes their bidding strategy and defines the volume they can bid for regulation service.
  2. 2a)
    Aggregator sends the volume they can bid for upward and downward regulation to TSO.
  3. 3a)
    TSO sends the MOL to DSO per ISP.
  4. 4a)
    DSO updates the boundary from the MOL and sends the new boundary to TSO.
  5. 5)
    TSO activates service in real-time.
The coordination between TSO and DSO starts from step 3a), after receiving the upper and lower boundary from the aggregator, TSO performs economic dispatch of flexibility to define the MOL and requests DSO for validation. Then, DSOs perform power flow analysis and inform TSO congestion state of the grid according to the traffic light concept mentioned in [20]. The green state means no congestion occurs. The yellow state informs the congestion happening in their network, DSO updates the upper and lower boundary of the flexibility by using linearized ACOPF and sends the new value to the TSO. Finally, the new boundary is sent to TSO.
II-B DSO-managed
Figure 2: DSO-managed coordination scheme.
Figure 2 illustrates the schemes where the DSO is prioritized in utilizing flexibility. The sequence of information exchange differs and is described below:
  1. 1.
    Aggregator optimal bidding strategy and define the volume they can bid for upward and downward regulation.
  2. 2b)
    Aggregators send the volume they can bid for upward and downward regulation to DSO.
  3. 3b)
    DSO sends the new boundary of the upward/downward volume that the aggregator can provide to TSO.
  4. 4)
    TSO defines the MOL and sends activation requests to the aggregator in real-time.
@@ II-B DSO-managed
The difference with the previous coordination schemes lies in step 2b, the aggregator will send their flexibility boundaries to DSO for validation instead of to the TSO. DSO updates the generation and demand of the grid with all flexibility and performs power flow analysis to check the congestion state in their network. Then DSO applies linearized ACOPF to determine the new boundaries of flexibility. Until no congestion occurs, DSO sends new boundaries to TSO. Finally, TSO defines the MOL based on the new boundaries.
@@
In the above processes, the aggregator performs its functions the day before, while the TSO and DSO exchange data during the offering of aFRR services. The service cost and benefit are determined when the market closes. Moreover, when providing services to the TSO, the aggregator deviates from the day-ahead schedule submitted to the BRP the day before [21]. As a result, the aggregator incurs a deviation fee payable to the BRP to compensate for the imbalance. This fee influences the aggregator’s decisions regarding how much flexibility to reserve. In addition to the deviation fee, the TSO’s minimum bid size presents another challenge for aggregators. To facilitate greater DER flexibility, the TSO should consider lowering the minimum bid size and increasing procurement times to reduce the uncertainty associated with forecast errors [22]. Therefore, this study assumes that TSO will use all resources provided by the aggregator.
@@ III Mathematical models
The mathematical representation of TSO, DSO and aggregator model will be discussed in the following subsections.
@@ III-A Aggregator model
The aggregator links prosumers and markets by determining the optimal bidding schedule to purchase electricity from the day-ahead market to sell to prosumers and to purchase services from prosumers to sell to other markets. Their objective is to maximize total benefit when providing services. The aggregator objective function and the relevant constraints for the whole day (96 steps) is present in Eq. 1 as below:
$F^{Agg}=\sum_{t}^{T}\sum_{i}^{N}E_{i,t}^{\uparrow}(\lambda_{t}^{\uparrow}-\lambda^{\text{brp}})+E_{i,t}^{\downarrow}(\lambda_{t}^{\downarrow}+\lambda^{\text{brp}})+E_{i,t}^{\text{da}}(\lambda_{t}^{\text{da}}-\lambda^{\text{p}})$ (1)
@@
$\sum_{t\in T}E_{i,t}^{\uparrow}=0\quad\forall i,t$ $\in[T_{\text{d},i},T_{\text{a},i}])$ (2)
$\sum_{t\in T}E_{i,t}^{\downarrow}=0\quad\forall i,t$ $\in[T_{\text{d},i},T_{\text{a},i}])$ (3)
$\sum_{t\in T}E_{i,t}^{\text{da}}=0\quad\forall i,t$ $\in[T_{\text{d},i},T_{\text{a},i}])$ (4)
$P^{\uparrow}_{i,t}=E^{\uparrow}_{i,t}/\Delta T$ $\quad\forall{t,i}$ (5)
$P^{\downarrow}_{i,t}=E^{\downarrow}_{i,t}/\Delta T$ $\quad\forall{t,i}$ (6)
$P^{da}_{i,t}=E^{da}_{i,t}/\Delta T$ $\quad\forall{t,i}$ (7)
@@
$E^{\text{EV}}_{i,t}=E^{\text{EV,Size}}_{i}$ $\quad\forall i,t=0$ (13)
$E^{\text{EV}}_{i,t}=E^{\text{EV,Size}}_{i}\quad\forall i,t=T_{d,i}$ $(\neq 0)$ (14)
$E^{\text{EV}}_{i,t}=E^{\text{EV,Size}}_{i}\quad\forall i,t=95$ $(\neq T_{\text{a}})$ (15)
@@
The volumes of flexibility used by each EV for upward and downward and the volume of energy bought from the day-ahead market are $E_{i}^{\uparrow}$, $E_{i}^{\downarrow}$, and $E_{i}^{\text{da}}$ accordingly. The set of $\lambda^{\uparrow}$, $\lambda^{\downarrow}$ and $\lambda^{\text{da}}$ are balancing price of flexibility for regulating upward, downward and the day-ahead price. $\lambda_{\text{brp}}$ is the contracted price between aggregator with BRP to pay for their deviation with the day-ahead schedule. $\lambda_{\text{p}}$ is the price aggregator sell electricity to the EV’s owner for charging their EV. And $N$ is the number of EVs considered.
Aggregator considers Eq. 2, Eq. 3 and Eq. 4 to make sure the EV will not interact with the electricity market when EVs are away from home, it means the total of energy provided for upward or downward regulation or buy from the DA market is zeros after the departure time $(T_{d})$ and before arrival time $(T_{a})$. In this study, EV is assumed to have only one leave session per day. Equations 5, 6 and 7 present the relationship between power $(P_{i})$ and energy $(E_{i})$ within one ISP, $\Delta T$ is 0.25 hour. Besides, power charge/discharge is limited to the min/max charging rate $(\underline{P_{i}^{\downarrow}},\overline{P_{i}^{\downarrow}})$ or min/max discharging rate $(\underline{P_{i}^{\uparrow}},\overline{P_{i}^{\uparrow}})$ by Eq. 8, 9 and 10. Constraint 11 ensures that the aggregator will provide multiple services with the ”In time” stacking type. This means that the EV will not provide up or downward service or buy from the DA market at the same time thanks to the use of the binary variables $u_{i,t},v_{i,t}$ and $w_{i,t}$. Moreover, the state of charge (SOC) of each EV is constrained by Eq. 12 means that EV will not charge over maximum capacity (100%) and will not discharge lower than 20% because deeper discharges can lead to accelerated wear and reduced lifespan of the battery.
At the beginning of the day, before leaving home and before midnight EV energy $(E^{\text{EV}}_{i,t})$ is initialized to be fully charged in Eq. 13, Eq. 14, Eq. 15. Equation 16 updates the EV’s energy by step when EV is at home, EV energy at step $t$ is equal to the energy from the previous step with the change from total energy charge and discharge. When EV leaves home, 17 energy is assumed to decrease gradually by the total energy for the trip ($E_{i}^{EV,trip}$) divided by the trip length ($TL$).

### v2

# Balancing service provision by EV aggregator in different TSO-DSO coordination schemes

@@ Abstract
The increasing penetration of Distributed Energy Resources (DERs) in the distribution system has led to the emergence of a new market actor - the aggregator. Aggregator serves as a facilitator, enabling flexibility asset owners to access various markets. Among them, electric vehicle (EV) aggregators are gaining more attention due to their expanding use and potential to provide services across multiple markets, particularly in the reserve market. Currently, the transmission system operator (TSO) indirectly utilizes the resources under the management of the distribution system operators (DSO), which can negatively impact the distribution grid. Conversely, adjustments from DSOs can impact service provision to TSO due to the shortage of TSO usage information. These factors highlight the importance of evaluating the service provision from aggregators under different TSO-DSO coordination schemes. This paper focuses on the provision of flexibility from electric vehicles (EVs) aggregators for balancing service in the TSO-DSO hybrid-managed scheme and compares it to the DSO-managed coordination schemes. The behavior of aggregators reacting to price fluctuations and TSO requests under different coordination schemes and simulation scenarios is thoroughly assessed. Additionally, their impact on the grid is analyzed through the DSO’s congestion management process and validated using data from a real part of the Dutch distribution network. Results find that the hybrid-managed coordination scheme gives more benefit to the aggregator than the DSO-managed scheme, and the price paid for the BRP for the deviation has a large impact on the aggregator’s behavior.
@@
TSO-DSO coordination, DERs, EVs aggregator, congestion, flexibility.
@@ I Introduction
Along with exploiting the potential of distributed energy resources (DERs), providing flexibility services with these sources attracts the attention of many organizations and research groups and has already been implemented in some European countries like France, Finland, Hungary, Estonia, Denmark, Belgium, Romania, etc [1]. The role of the aggregator - a market participant responsible for bundling flexibility from multiple small customers and producers into a portfolio and offering the combined capacity in other markets[2] - is becoming increasingly important. Many research studies focus on optimal biding strategies[3, 4, 5, 6, 7, 8], the provision of services to multiple markets (e.g., spot market, reserve,[3, 9, 10, 11, 12] congestion management services[7] or balancing portfolios[3]) and the utilization of diverse appliances including electric vehicles (EVs), battery energy storage system (BESS), heat pumps, thermal-controlled load or their combinations.
@@
Recently, EV aggregators have been receiving increasing attention due to their expanding use, driven by environmental factors and government incentives. In 2023, the Netherlands was the sixth-largest EV market in Europe with over 300 thousand battery electric vehicles registered and leads in EV infrastructure [13]. Research on the Dutch situation shows the potential of EV aggregators to provide service in the reserve market[10]. EV fleets also demonstrate potential to provide flexibility to the internal balancing portfolio[3], and congestion management [7]. The above studies show that EV aggregators have great prospective to provide flexibility services and bring benefits not only for aggregators but also for prosumers and system operators. Therefore, the EV aggregator is investigated in this research to consider its behavior on different market price fluctuations.
@@
Moreover, previous research has primarily focused on the aggregator business model when providing service to the transmission system operator (TSO), with limited attention given to the impact on the grid under the management of the distribution system operator (DSO). While an effective approach to leveraging DERs flexibility is enhancing TSO-DSO coordination [14]. In recent years, this coordination has become a prominent research trend, attracting significant interest from the academic community. Arthur et al [15] classify the existing coordination into three main schemes (1) TSO-managed; (2) DSO-managed; and (3) TSO-DSO Hybrid. In practice, the DER service provider contracts directly with the TSO in the first model. TSO is the single buyer and is prioritized in accessing DER flexibility and performs economic dispatch of DERs, while the DSO is restricted from accessing resources connected to their network. With less interaction, the implementation of this scheme is straightforward in a centralized manner. However, the distribution network constraints are not considered due to the lack of information. This scheme also poses great computational and modeling challenges to TSO, and a large amount of operational data needs to be shared. Conversely, the DSO-managed model grants the DSO a more prominent role and greater responsibility in optimizing the use of flexibility. The DSO is given priority to aggregate and employ the DER flexibility for managing congestion in their grids, then send the remaining flexibility to the central market. With increasing RES, it becomes more difficult to implement in a decentralized way via DSOs. Moreover, this scheme has the potential DSO conflicts of interest and transfers the computational challenges to DSO. Among these 3 schemes, the Hybrid TSO-DSO managed model appears to be the most balanced option as TSO and DSO manage their own market and allow TSO to use service with DSO’s consent. The aggregator sends their bid to TSO, then TSO asks DSO to validate the bid. This coordination aligns well with the prevailing circumstances in many countries and offers distinct advantages over others by empowering DSOs role and fostering greater social welfare. Therefore, the DSO-TSO hybrid-managed model is one of the schemes selected for investigation in this research.
@@
Considering the provision of DERs flexibility from the aggregator within the TSO-DSO coordination schemes offers a broader perspective by highlighting the risk of violating grid constraints when procuring flexibility for the TSO. Therefore, this research constructs a comprehensive model in which an EV aggregator provides multiple services under different TSO-DSO coordination schemes. To the best of the authors’s knowledge, this topic has not been widely explored in the literature. This paper shares some similarities with the 3 papers [16, 17, 18] in using flexibility from the distribution grid via the aggregator while accounting for the TSO-DSO coordination. However, unlike [16], this paper models the details of the EV aggregator, improves the DSO model with linearized ACOPF and compares the flexibility provision in 2 different coordination schemes. In [17], the authors show that 16% and nearly 52% of the operational cost is reduced for TSO and DSO, respectively. While the aggregator experienced a profit loss of approximately 54% under the DSO-managed model compared to the case without coordination, the study did not explore how the outcomes would differ if the TSO utilized flexibility from the aggregator in the balancing market. Compared with [18], this paper models the EV aggregator that considers individual EV to ensure EV energy within the acceptable limit range while providing vehicle-to-grid service and considering the TSO-DSO hybrid-managed model that advances over the DSO-managed coordination scheme.
@@
II TSO-DSO coordination
Coordination between TSOs and DSOs can take place in various stages, including pre-qualification, offering, and real-time activation of services. In the context of the Dutch balancing service, coordination is primarily limited to the pre-qualification phase, where equipment specifications are reviewed before granting grid connection approval. However, no interaction occurs during the offering and activation stages, which are conducted in real-time or near real-time In this study, coordination in the offering process for automatic frequency restoration reserve (aFRR) provision is proposed with two different schemes: the DSO-managed model and the TSO-DSO hybrid-managed model. The TSO-managed model places a significant computational burden on the TSO and limits the DSO’s ability to flexibly utilize services from resources connected to its own network. This restriction can lead to reduced efficiency and fairness in the facilitation of DER flexibility services. As a result, this type of coordination is considered beyond the scope of this study.
II-A TSO-DSO hybrid-managed coordination scheme
Figure 1: a) TSO-DSO hybrid-managed, and b) DSO-managed coordination schemes.
The first scheme is illustrated in Fig. 1a on the left, which consists of 5 main processes. In which, (1) the aggregator aggregates all the resources from their prosumers, determines the optimal bidding strategy and sends it to the balance responsible party (BRP) before the energy market gate closing time. (2a) The aggregator also defines the portion of flexibility to be offered as balancing services and sends this information to the TSO. (3a) The TSO conducts the economic dispatch and compiles the merit order list (MOL) (a list of offers sorted by price from low to high) for the next two imbalance settlement period (ISP) and forwards it to the DSO. (4a) Upon receiving the MOL, the DSO performs the validation process. This involves eliminating invalid bids/offers by perform power flow analysis to determine the grid congestion, following the traffic light concept described in [19]. The green state indicates that no congestion occurs, while the yellow state signals that congestion is happening in their network. DSO updates the upper and lower boundary of the flexibility by using linearized ACOPF and communicates the validated result back to the TSO. (5) Finally, TSO activates the service in real time.
II-B DSO-managed coordination scheme
Fig. 1b on the right illustrates the schemes where the DSO is prioritized in utilizing flexibility. The sequence of information exchange is structured into four main processes. (1) The aggregator optimizes its bidding strategy and defines the volume available for upward and downward regulation and (2b) These bids and offers are then submitted to the DSO for validation. Upon validation, DSO sends the volume they use to resolve congestion to the aggregator (3b) and the new boundary of the upward/downward volume that the aggregator can provide to TSO (4b). Based on the new boundaries, TSO constructs the MOL and sends activation requests to the aggregator (5).
@@ II-B DSO-managed coordination scheme
The key difference between this coordination scheme with the previous one lies in steps (2b) and (3b), where the aggregator sends its flexibility boundaries to DSO for validation, rather than directly to the TSO, then DSO sends back the volume they use to resolve congestion to the aggregator. DSO updates the grid’s generation and demand profiles by incorporating all available flexibility and performs power flow analysis to check the congestion state in the network. Subsequently, the DSO applies linearized ACOPF to determine updated boundaries of flexibility. This process is repeated until no congestion is detected, the DSO transmits new boundaries to the TSO. Finally, TSO defines the MOL based on the new boundaries. Unlike the DSO-managed model in [15], the TSO will send its dispatch command to the DER aggregator without DSO acting as an intermediate, due to the dispatched bid is assumed unchanged through the DSO.
@@
In the above processes, the aggregator performs its functions the day before, while the TSO and DSO communicate data during the offering of the aFRR services. The service cost and benefit are determined when the market closes. Moreover, when providing services to the TSO, the aggregator deviates from the day-ahead schedule submitted to the BRP in the day before [20]. As a result, the aggregator incurs a deviation fee payable to the BRP to compensate for their imbalance. This fee influences the aggregator’s decisions regarding how much flexibility to reserve. In addition to the deviation fee, the TSO’s minimum bid size poses another barrier for aggregators. To enable greater DER flexibility, the TSO should consider lowering the minimum bid size and increasing procurement times, thereby helping to mitigate uncertainty stemming from forecast errors [21]. Therefore, this study assumes that TSO will use all resources provided by the aggregator.
@@ III Mathematical models
The mathematical formulations of the TSO, DSO and aggregator models will be discussed in the following subsections.
@@ III-A Aggregator model
The aggregator links prosumers and markets by determining the optimal bidding schedule to purchase electricity from the day-ahead market to supply prosumers and procuring services from prosumers to offer in other markets. The aggregator’s objective is to maximize total benefit when providing services. The objective function and the associated constraints, defined over the whole day (96 steps), are presented below:
$max\sum_{t}^{T}\sum_{i}^{N}E_{i,t}^{\uparrow}(\lambda_{t}^{\uparrow}-\lambda^{\text{brp}})+E_{i,t}^{\downarrow}(\lambda_{t}^{\downarrow}+\lambda^{\text{brp}})+E_{i,t}^{\text{da}}(\lambda_{t}^{\text{da}}-\lambda^{\text{p}})\$ (1)
@@
$P^{\uparrow/\downarrow/da}_{i,t}=E^{\uparrow/\downarrow/da}_{i,t}/\Delta T$ $\quad\forall{t,i}$ (2)
@@
$E^{\text{EV}}_{i,t}=E^{\text{EV,Size}}_{i}\quad\forall i,t=0/T_{d,i}/95$ (8)
@@
The volumes of flexibility of each EV for upward and downward and the volume of energy bought from the day-ahead market are $E_{i}^{\uparrow}$, $E_{i}^{\downarrow}$, and $E_{i}^{\text{da}}$ accordingly. The set of $\lambda^{\uparrow}$, $\lambda^{\downarrow}$ and $\lambda^{\text{da}}$ are balancing prices of flexibility for regulating upward, downward and the day-ahead price, respectively. $\lambda_{\text{brp}}$ represents the contracted price that the aggregator pays to the BRP for their deviation from the day-ahead schedule. $\lambda_{\text{p}}$ denotes the price at which the aggregator sells electricity to the EV’s owner for charging. And $N$ refers to the number of EVs considered, while T indicates the number of time step considered in the next day, which is set to 96 in this study. Each EV is assumed to have a single departure session per day and to charge exclusively at home. This implies that there is no interaction with the electricity market after the departure time $(T_{d})$ and before the arrival time $(T_{a})$. Eq. 2 presents the relationship between power $(P_{i})$ and energy $(E_{i})$ within one ISP, $\Delta T$ is 0.25 hour. Besides, power charge/discharge is limited to the min/max charging rate $(\underline{P_{i}^{\downarrow}},\overline{P_{i}^{\downarrow}})$ or min/max discharging rate $(\underline{P_{i}^{\uparrow}},\overline{P_{i}^{\uparrow}})$ by Eq. 3, 4 and 5. Constraint 6 ensures that the aggregator will provide multiple services with the ”In time” stacking type. This implies that the EV is neither provides upward nor downward service nor buys from the DA market at the same time step. This exclusivity is ensured through the use of the binary variables $u_{i,t},v_{i,t}$, and $w_{i,t}$. Moreover, the state of charge (SOC) of each EV is constrained by Eq. 7, ensuring that SOC does not exceed the maximum capacity (100%) and does not fall below 20%. This lower bound prevents deeper discharges, which can lead to accelerated wear and reduced lifespan of the battery. At the beginning of the day, at the leaving moment, and at midnight, EV energy $(E^{\text{EV}}_{i,t})$ is initialized to be fully charged in Eq. 8. Equation 9 updates the EV’s energy by step when the EV is at home, EV energy at step $t$ is equal to the energy from the previous step, incorporating the change from total energy charge and discharge. When EV leaves home, energy is assumed to decrease gradually by the total energy for the trip ($E_{i}^{EV,trip}$) divided by the trip length ($TL$) via Eq. 10.